# Audio-Agent: Leveraging LLMs For Audio Generation, Editing and Composition

### v1

@@ Abstract
We introduce Audio-Agent, a multimodal framework for audio generation, editing and composition based on text or video inputs. Conventional approaches for text-to-audio (TTA) tasks often make single-pass inferences from text descriptions. While straightforward, this design struggles to produce high-quality audio when given complex text conditions. In our method, we utilize a pre-trained TTA diffusion network as the audio generation agent to work in tandem with GPT-4, which decomposes the text condition into atomic, specific instructions, and calls the agent for audio generation. Consequently, Audio-Agent generates high-quality audio that is closely aligned with the provided text or video while also supporting variable-length generation. For video-to-audio (VTA) tasks, most existing methods require training a timestamp detector to synchronize video events with generated audio, a process that can be tedious and time-consuming. We propose a simpler approach by fine-tuning a pre-trained Large Language Model (LLM), e.g., Gemma2-2B-it, to obtain both semantic and temporal conditions to bridge video and audio modality. Thus our framework provides a comprehensive solution for both TTA and VTA tasks without substantial computational overhead in training.
@@ 1 Introduction
Multimodal deep generative models have gained increasing attention these years. Essentially, the models are trained to perform tasks based on different kinds of input called modalities, mimicking how humans make decisions from different kinds of senses such as vision and smell Suzuki & Matsuo (2022). Compared to other generation tasks such as image generation or contextual understanding, audio generation is less intuitive as it is harder to precisely measure the generated sound quality using human ears. Additionally, previous works mainly focus on generating music-related audio, which is more structured compared to naturally occurring audio Copet et al. (2024); Melechovsky et al. (2023). Some recent works have focused on generating visually guided open-domain audio clips Chen et al. (2020); Zhou et al. (2018).
Recent researches on audio generation are mainly focused on text-to-audio generation (TTA) and video-to-audio generation (VTA). For TTA task Xue et al. (2024); Kreuk et al. (2022), current datasets lack high-quality text-audio pairs. Existing datasets such as AudioCaps Kim et al. (2019) or Clotho Drossos et al. (2020) usually contain multiple event descriptions mixed into one single sentence without fine-grained details and object bindings. This complicates training, particularly when handling long continuous signals with complex text conditions Huang et al. (2023). We define complex text conditions as long event descriptions containing a series of events without explicitly describing the sound, such as “A man enters his car and drives away”. While previously not fully studied, this type of condition is more realistic as it does not require any detailed specification in terms of the characteristics of the audio result, offering more flexibility to the user and producer for areas such as movie dubbing and musical composition. If we train these models from scratch, it often demands extensive computational resources Liu et al. (2024); Ghosal et al. (2023).
The VTA task, or conditional Foley generation, remains unexplored until recently Wang et al. (2024); Zhang et al. (2024b). One main challenge is that video clips typically contain excessive visual information not always relevant to audio generation. Moreover, synchronization is hard between video and audio output, with recent solutions such as temporal masking Xie et al. (2024) proving inadequate for complex scenarios. Due to efficiency considerations, current methods often encode video features by extracting a few random frames Xie et al. (2024); Dong et al. (2023), which hinders learning temporal information. Bridging the modality gap Liang et al. (2022) between video and audio thus becomes the key to solving the problem.
While achieving state-of-the-art results, conventional approaches often perform inference in a single pass based on a given text description. This approach struggles to produce high-quality audio when faced with complex or lengthy text conditions. In this paper, we introduce Audio-Agent, which breaks down intricate user inputs using GPT-4 into multiple generation steps. Each step includes a description along with start and end times to effectively guide the audio generation process. Our framework integrates two key tasks: Text-to-Audio (TTA) and Video-to-Audio (VTA). We leverage a pre-trained TTA diffusion model, Auffusion Xue et al. (2024), with essential adaptations, serving as the backbone for our generation process. In the TTA task, Auffusion focuses solely on generating simple, atomic text inputs. Our framework supports audio generation, editing, and composition, as illustrated in Figure 1. For the VTA task, we recognize that models such as GPT-4 and other large language models lack sufficient temporal understanding of video clips. To address this problem, we employ moderate fine-tuning to align the two modalities. We utilize the smaller Gemma2-2B-it model, which has 2 billion parameters, and fine-tune an adapter and a projection layer to convert visual inputs into semantic tokens. We then implement cross-attention guidance between the diffusion layers of Auffusion. This approach eliminates the need for additional training on a temporal detector, as the semantic tokens inherently contain time-aligned information.
The summary of our contributions is as follows: 1) we propose Audio-Agent which utilizes a pre-trained diffusion model as a generation agent, for both TTA and VTA tasks; 2) For TTA, Audio-Agent can handle complex text input, which is broken down into simple and atomic generation conditions for the diffusion model to make inference on; 3) For VTA, we fine-tune an open-source LLM (Gemma2-2B-it) to bridge the modality gap between video and audio modalities to align the underlying semantic and temporal information. Through extensive evaluation, our work demonstrates on-par results compared to the state-of-the-art task-specific models trained from scratch, while capable of producing high-quality audio given long and complex textual input. We hope our work can motivate more relevant works on multi-event long-condition TTA generation, which to our knowledge has not yet been fully explored despite its high potential in various content generations where high-quality audio is essential.
@@
Figure 1: Example showing Audio-Agent’s ability to generate, compose and edit multiple audio descriptions together. (A): Multi-turn editing; (B): Generation based on long description; (C): Multiple audio descriptions composition
@@ 2 Related Work
LLM-based Agent Method Recent progress in large language models has enabled relevant research on making LLM a brain or controller for the agent on performing various tasks, such as robot task planning and execution Driess et al. (2023) or software development Rawles et al. (2024); Yang et al. (2023). LLM demonstrates the capacity of zero-shot or few-shot generalization, making task transfer possible without significant change of its parameters Xi et al. (2023). In our work, we harness the action-planning ability of LLM. Upon receiving the text condition from the user, LLM generates a plan with detailed steps on how to call the diffusion model which serves as a generation agent. By dividing the task into simpler sub-tasks, we can ensure the generation quality with fine-grained event control for TTA generation.
@@
Diffusion-based Audio Generation AudioLDM Liu et al. (2024) is among the pioneering works that introduce the latent diffusion method to audio generation. Subsequent works such as Tango Ghosal et al. (2023) and Auffusion Xue et al. (2024) use pre-trained LLM such as Flan-T5 for text encoding, which has been widely adopted. We notice that this method can be seamlessly adapted to VTA tasks when we can find a similarly effective way of utilizing LLM for encoding the visual content. For the TTA task, we choose Auffusion as our generation agent due to its outstanding performance on fine-grained alignment between text and audio.
@@
Coarse-to-fine Audio Generation Current works such as AudioLM Borsos et al. (2023), VALL-E Wang et al. (2023) and MusicLM Agostinelli et al. (2023) use multiple codebooks and Residual Vector Quantization (RVQ) Défossez et al. (2022) to create diverse audio representations. In AudioLM, the model first predicts semantic tokens that capture crucial information for overall audio quality, such as rhythm and intonation, while subsequent layers add details to enhance the richness of the generated sound. However, these discrete designs suffer from generation quality compared to their continuous-valued counterparts. Moreover, the model has to perform prediction over multi-layers, which inevitably increases computational demands for both training and inference Meng et al. (2024). In our case for the VTA task, we fine-tune an LLM to predict an intermediate discrete representation as semantic tokens using a language modeling approach. The discrete semantic tokens then serve as a condition for the diffusion model to generate continuous predictions. In this way, our method simplifies the generation procedure while maintaining the advantages of audio generation using the language modeling approach.
@@ 3 Method
3.1 Preliminaries
Audio Latent Diffusion Model Recent research adapted the successful latent diffusion models from the image domain to the audio domain. A typical audio latent diffusion model such as Auffussion first converts the audio wave into mel spectrogram, followed by VAE encoding into the relevant latent space. Inference is the reverse process, where the predicted latent is decoded by VAE and then converted back from mel spectrogram into audio wave through a vocoder such as HiFi-GAN Kong et al. (2020). The latent diffusion process can be regarded as the same as the standard latent diffusion model on image generation Rombach et al. (2022).
Semantic token AudioLM Borsos et al. (2023) was among the first to propose a two-stage method for speech synthesis. In their method, the semantic tokens are derived from representations produced by an intermediate layer of w2v-BERT Chung et al. (2021). We choose an open-sourced HuBERT Hsu et al. (2021) model to produce the semantic representation, since HuBERT can model long-term temporal structure in a generative framework. Although only the smallest Hubert model has its quantizer released and open-sourced, we found that the released small model is already enough to assist the diffusion model in generating high-quality and temporally aligned predictions.
@@ 3.2 GPT-4 as an action planner for TTA task
Given a long, complex text condition, we ask GPT-4 to decompose the description into simple and atomic generation steps. GPT-4 has the freedom to decide how many steps to generate. We additionally restrict GPT-4 to keep the minimum number of necessary generation steps. This step instruction produces a good balance avoiding either extreme of being too abstract or too specific with unnecessary details. We also inform GPT-4 that the user may revise the text requirement in subsequent conversations so that our framework can perform multi-turn conversational generation. The output of GPT-4 consists of a JSON file, which contains a series of function calls of the agent model with text description provided. In addition, to support variable length generation and multi-event generation, GPT-4 also provides the start time and end time for each call which can overlap with each other. After obtaining the generation result for each step, we add waveforms together based on their time range. See Appendix A.1 for a prompt example.
Table 1: Comparison of functionalities between recent audio generation framework. For AudioLDM2 and Auffusion half check marks are assigned because the corresponding model was trained only on 10 seconds of audio clips. In theory, it also supports long audio generation, but the quality is not assured, see Figure 5
. Method VTA generation TTA generation Multi-turn editing Composition Long complex generation Diff-Foley FoleyCrafter AudioLDM2 Auffusion Ours
@@ 3.3 Audio Tokenizer and Video Tokenizer
Following Kharitonov et al. (2021), we utilize the 9th layer of the Hubert-Base model to derive the semantic tokens. The quantizer of Hubert-Base contains 500 centroids. Given an audio clip as ground truth, Hubert acts as an audio tokenizer that applies K-mean clustering and converts the audio into discrete semantic tokens, where each token has a value ranging from 0 to 499 to represent the respective centroids. Hubert-Base has a frame rate of 50Hz, thus a 10-second audio will result in 500 semantic tokens.
To efficiently capture both visual and temporal information while compressing the video data, we employ CLIP as a frame-wise feature tokenizer. CLIP is compatible with arbitrary frame sampling strategies, enabling a more flexible frame-to-video feature aggregation scheme as noted by Cheng et al. (2024). We pool the information within each frame to reduce the sequence size, resulting in a vector $f^{r}$ of size $N\times D$, where $N$ is the number of frames and $D$ is the CLIP hidden size. We set the frame rate to 21.5 Hz and use CLIP ViT-L/14 by default.
Inter-frame information is crucial for the model to achieve temporal alignment. Previous methods Iashin & Rahtu (2021); Du et al. (2023) require extracting both RGB and optical flow information within and across frames. In our design, we add a temporal connector after obtaining frame-wise features. The temporal connector consists of a 1D convolution block and a projection layer. The convolution block aggregates the inter-frame features together while preserving the temporal order. The projection layer projects the features into LLM’s embedding space.
@@ 3.4 LLM for semantic token generation on VTA task
Semantic tokens allow us to represent continuous audio information in discrete semantic form. We denote the continuous audio ground truth as $a\in\mathbb{R}^{C\times L}$, where $C$ is the number of channels and $L$ is the time of the audio clip times sample rate. The Hubert audio tokenizer applies the K-means algorithm to convert the representation into LLM-aware acoustic tokens. Specifically, we obtain the indices $s\in\{0,...,499\}^{N}$ from the audio by comparing it with the encoded audio with centroids, and $N$ is the sequence length.
@@
In our experiments, we use Gemma2-2B-it Team et al. (2024), a lightweight open-source LLM developed by Google, which is claimed to have comparable performance to a much larger variant Gemma-2-9B. We use Low-Rank Adaptor (LoRA) Hu et al. (2021) to finetune Gemma to make it understand vision/text conditions and generate audio tokens.
Figure 4: A demo example showing Audio-Agent’s conversation ability: First turn: Audio Generation; second turn: Audio Insertion; third Turn: Audio Editing; last turn: Audio Composition with high-level semantic instructions. Audio-Agent can choose to respond based on previous turns or make independent generations.

### v2

@@ Abstract
We introduce Audio-Agent, a multimodal framework for audio generation, editing and composition based on text or video inputs. Conventional approaches for text-to-audio (TTA) tasks often make single-pass inferences from text descriptions. While straightforward, this design struggles to produce high-quality audio when given complex text conditions. In our method, we utilize a pre-trained TTA diffusion network as the audio generation agent to work in tandem with GPT-4, which decomposes the text condition into atomic, specific instructions and calls the agent for audio generation. In doing so, Audio-Agent can generate high-quality audio that is closely aligned with the provided text or video exhibiting complex and multiple events, while supporting variable-length and variable-volume generation. For video-to-audio (VTA) tasks, most existing methods require training a timestamp detector to synchronize video events with the generated audio, a process that can be tedious and time-consuming. Instead, we propose a simpler approach by fine-tuning a pre-trained Large Language Model (LLM), e.g., Gemma2-2B-it, to obtain both semantic and temporal conditions that bridge the video and audio modality. Consequently, our framework contributes a comprehensive solution for both TTA and VTA tasks without substantial computational overhead in training.
@@ 1 Introduction
Multimodal deep generative models have gained increasing attention these years. These models are trained to perform tasks based on different kinds of input called modalities, mimicking how humans make decisions from different sensory inputs such as vision and smell [38]. Compared to other generation tasks such as image generation or contextual understanding, audio generation is less intuitive as it is harder to precisely measure the generated sound quality using human ears. Moreover, previous works mainly focus on generating music-related audio, which is more structured compared to naturally occurring audio [7, 33]. Some recent works have focused on generating visually guided open-domain audio clips [4, 53].
@@
Figure 1: Example showing Audio-Agent’s ability to generate, compose and edit multiple audio descriptions together. (A): Multi-turn editing; (B): Generation based on long description; (C): Multiple audio descriptions composition.
Recent research on audio generation has mainly focused on text-to-audio generation (TTA) and video-to-audio generation (VTA). For TTA [48, 24], current datasets lack high-quality text-audio pairs. Existing datasets such as AudioCaps [22] or Clotho [11] usually contain multiple event descriptions mixed into a single sentence without fine-grained details and object bindings. This complicates training, particularly when handling long continuous signals with complex text conditions [17]. We define complex text conditions as long event descriptions containing a series of events without explicitly describing the sound, such as “A man enters his car and drives away”. While previously not fully studied, this type of condition is more realistic, as it does not require any detailed specification in terms of the characteristics of the audio result, offering more flexibility to users and producers in movie dubbing and musical composition.
The VTA task, or conditional Foley generation, has remained relatively unexplored until recently [42, 52]. One main challenge is that video clips typically contain excessive visual information not always relevant to audio generation. Moreover, synchronization is hard between video and audio output, with recent solutions such as temporal masking [46] showing still much room for improvement to handle complex scenarios. Due to efficiency considerations, current methods often encode video features by extracting a few random frames [46, 9], which hinders learning temporal information. Bridging the modality gap [27] between video and audio thus becomes the key to solving the problem.
While achieving state-of-the-art results, conventional approaches perform inference in a single pass based on a given text description. This approach struggles to produce high-quality audio when given complex or lengthy text conditions as input. In this paper, we introduce Audio-Agent, which employs GPT-4 to delineate intricate user inputs using GPT-4 into multiple generation steps. Each step includes a description along with start and end times and optionally volume control to effectively guide the audio generation process. Our general framework integrates two key tasks: Text-to-Audio (TTA) and Video-to-Audio (VTA).
First, we leverage a pre-trained TTA diffusion model, Auffusion [48] with essential adaptations, serving as the backbone for our generation process. In the TTA task, Auffusion focuses solely on generating simple, atomic text inputs. Our framework supports audio generation, editing, and composition, as illustrated in Figure 1.
Second, for the VTA task, we recognize that models such as GPT-4 and other large language models lack sufficient temporal understanding on video clips. To address this problem, we employ moderate fine-tuning to align the two modalities. We utilize the smaller Gemma2-2B-it model, which has 2 billion parameters, and fine-tune an adapter and a projection layer to convert visual inputs into semantic tokens. We then implement cross-attention guidance between the diffusion layers of Auffusion. This approach eliminates the need for additional training on a temporal detector, as the semantic tokens inherently contain time-aligned information. Our contributions are summarized as:
  1. 1.
    we propose Audio-Agent, which utilizes a pre-trained diffusion model as a generation agent, for both TTA and VTA tasks;
  2. 2.
    for TTA, Audio-Agent can handle complex text input, which is decomposed into simpler and atomic generation conditions on which the diffusion model can effectively make inference;
  3. 3.
    for VTA, we fine-tune an open-source LLM (Gemma2-2B-it) to bridge the modality gap between video and audio modalities to align the underlying semantic and temporal information.
Through extensive evaluation, our work demonstrates on-par results compared to the state-of-the-art task-specific models trained from scratch, while capable of producing high-quality audio given long and complex textual input. We hope our work can motivate more relevant works on multi-event long-condition TTA generation, which to our knowledge has not yet been adequartely explored despite its high potential in various content generations where high-quality audio is essential.
@@ 2 Related Work
LLM-based Agent Method Recent progress in large language models has enabled relevant research on making LLM a brain or controller for the agent on performing various tasks, such as robot task planning and execution [10] or software development [35, 49]. LLM demonstrates the capacity of zero-shot or few-shot generalization, making task transfer possible without significant change of its parameters [45]. In our work, we harness the action-planning ability of LLM. Upon receiving the text condition from the user, LLM generates a plan with detailed steps on how to call the diffusion model which serves as a generation agent. By dividing the task into simpler sub-tasks, we can ensure the generation quality with fine-grained event control for TTA generation.
Table 1: Comparison of functionalities between recent audio generation framework. For AudioLDM2 and Auffusion half check marks are assigned because the corresponding model was trained only on 10 seconds of audio clips. In theory, it also supports long audio generation, but the quality is not assured, see Figure 5.
Method VTA generation TTA generation
Multi-turn editing Composition Long complex generation
Diff-Foley
FoleyCrafter
AudioLDM2
Auffusion
Make-An-Audio 2
AudioGPT
WavJourney
WavCraft
Ours
@@
Diffusion-based Audio Generation AudioLDM [28] is among the pioneering works that introduce the latent diffusion method to audio generation. Subsequent works such as Tango [13] Auffusion [48], and AudioLDM 2 [29] use pre-trained LLM such as Flan-T5 for text encoding, which has been widely adopted. We notice that this method can be seamlessly adapted to VTA tasks when a similarly effective way can be derived for utilizing LLM in encoding visual content. For the TTA task, we choose Auffusion as our generation agent due to its outstanding performance on fine-grained alignment between text and audio.
@@
Coarse-to-fine Audio Generation Current works such as AudioLM [2], VALL-E [41] and MusicLM [1] use multiple codebooks and Residual Vector Quantization (RVQ) [8] to create diverse audio representations. In AudioLM, the model first predicts semantic tokens that capture crucial information for overall audio quality, such as rhythm and intonation, while subsequent layers add details to enhance the richness of the generated sound. However, these discrete designs suffer from generation quality compared to their continuous-valued counterparts. Moreover, the model has to perform prediction over multi-layers, which inevitably increases computational demands for both training and inference [34].
TTA and VTA Make-an-audio 2 [17] is one of the first methods capable of taking complex textual descriptions for audio generation. However, this method primarily focuses on the ordering of the generated audio, lacking the ability to further edit the generated audio through multi-turn conversations. Furthermore, [17] relies on extensive training with LLM-based data augmentation and is limited to TTA generation. Subsequent works, such as AudioGPT [18], WavJourney [30], and WavCraft [26], introduce specialized audio generators for compositional audio. However, these models still rely on off-the-shelf generators and do not fully leverage GPT-4’s potential. They are also limited to text-based descriptions, and the generated audio often fails to meet user-specified length requirements, as the complexity of GPT-4’s generation process exceeds their capabilities.
For VTA, recent works such as FoleyGen [32] and V2meow [37] employ transformers to model the audio generation process. However, as noted in [25], this autoregressive design can lead to suboptimal performance compared to continuous generation methods. In particular, errors tend to accumulate during inference, which can degrade the model’s overall performance.
@@ 3 Method
3.1 Preliminaries
Audio Latent Diffusion Model Recent research adapted the successful latent diffusion models from the image domain to the audio domain. A typical audio latent diffusion model such as Auffussion first converts the audio wave into mel spectrogram, followed by VAE encoding into the relevant latent space. Inference is the reverse process, where the predicted latent is decoded by VAE and then converted back from mel spectrogram into audio wave through a vocoder such as HiFi-GAN [23]. The latent diffusion process can be regarded as the same as the standard latent diffusion model on image generation [36].
Semantic token AudioLM [2] was among the first to propose a two-stage method for speech synthesis. In their method, the semantic tokens are derived from representations produced by an intermediate layer of w2v-BERT [6]. We choose an open-sourced HuBERT [15] model to produce the semantic representation, since HuBERT can model long-term temporal structure in a generative framework. Codec models such as w2v-BERT or HuBERT are originally pre-trained on speech-heavy datasets, still they exhibit the potential to be extended in other fields like music generation, as verified in [1]. Although only the smallest HuBERT model has its quantizer released and open-sourced, we found that the released small model is already sufficient to assist the diffusion model in generating high-quality and temporally aligned predictions.
@@ 3.2 GPT-4 as an action planner for TTA task
Given a long, complex text condition, we ask GPT-4 to decompose the description into simple and atomic generation steps. While GPT-4 has freedom to decide the number of steps to generate, we restrict GPT-4 to keep the minimum number of necessary generation steps. This step instruction produces a good balance avoiding either extreme of being too abstract or too specific with unnecessary details. We also inform GPT-4 that the user may revise the text requirement in subsequent conversations, so that our framework can perform multi-turn conversational generation.
The output of GPT-4 consists of a JSON file, which contains a series of function calls to the agent model with the text description provided. In addition, to support variable length generation and multi-event generation, GPT-4 also provides the start time and end time for each call which can overlap with each other. After obtaining the generation result for each step, we add the pertinent waveforms together based on their time range. During our experiments, we found that sometimes if the user provides compositional textual descriptions, such as combining lightning and people shouting events together, the overlapping region of the final audio may experience volume explosion. Thus, in one of our design variations, we further ask GPT-4 to control the audio volume by specifying the volume for each generation call in dB, following LUFS standard. See the supplementary material for prompt examples.
@@ 3.3 Audio Tokenizer and Video Tokenizer
Following [21], we utilize the 9th layer of the HuBERT-Base model to derive the semantic tokens. The quantizer of HuBERT-Base contains 500 centroids. During the training process of the VTA task, we aim to fine-tune the LLM to predict the correct semantic tokens indices from the video input. To obtain the ground truth semantic tokens indices, we first extract the ground truth audio clips from the video. We then use HuBERT acting as an audio tokenizer that applies the K-mean clustering algorithm and converts the audio into the target discrete semantic tokens indices, where each index has a value ranging from 0 to 499 to represent the respective centroids. During inference, the semantic token is constructed by fetching the corresponding centroid. HuBERT-Base has a frame rate of 50Hz, thus a 10-second audio will produce 500 semantic tokens. The HuBERT model and LLM act as a bridge that connects the video modality and audio modality. Compared with previous methods that directly use a projection layer for alignment, the semantic token serves as an intermediate representation between visual and audio modality, facilitating the conditional audio generation process.
To efficiently capture both visual and temporal information while compressing the video data, we employ CLIP as a frame-wise feature tokenizer. CLIP is compatible with any frame sampling strategies, enabling a flexible frame-to-video feature aggregation scheme as noted in [5]. We pool the information within each frame to reduce the sequence size, resulting in a vector $f^{r}$ of size $N\times D$, where $N$ is the number of frames and $D$ is the CLIP hidden size. We set the frame rate to 21.5 Hz and use CLIP ViT-L/14 by default.
Inter-frame information is crucial for the model to achieve temporal alignment. Previous methods [19, 12] require extracting both RGB and optical flow information within and across frames. In our design, we add a temporal connector after obtaining frame-wise features. The temporal connector consists of a 1D convolution block and a projection layer. The convolution block aggregates the inter-frame features together while preserving the temporal order. The projection layer projects the features into LLM’s embedding space.
@@ 3.4 LLM for semantic token generation on VTA task
Semantic tokens allow us to represent continuous audio information in discrete semantic form. We denote the continuous audio ground truth extracted from the video input as $a\in\mathbb{R}^{C\times L}$, where $C$ is the number of channels and $L$ is the time of the audio clip times sample rate. The HuBERT audio tokenizer applies the K-means algorithm to convert the representation into discrete indices. Specifically, we obtain the indices $s\in\{0,...,499\}^{N}$ from the audio by comparing it with the encoded audio with centroids, and $N$ is the sequence length.
@@
In our experiments, we use Gemma2-2B-it [39], a lightweight open-source LLM developed by Google, which is claimed to have comparable performance to a much larger variant Gemma-2-9B. We use Low-Rank Adaptor (LoRA) [16] to finetune Gemma to make it understand vision/text conditions and generate audio tokens.
Figure 4: A demo example showing Audio-Agent’s conversation ability: First turn: Audio Generation; second turn: Audio Insertion; third Turn: Audio Editing; last turn: Audio Composition with high-level semantic instructions. Audio-Agent can choose to respond based on previous turns or make independent generations. The corresponding audio files can be found in the supplementary materials.